10g base 4x+log base 4(x+1)=log base 4(20)
Answer by stanbon(75887)

1og base4(x) + log base4(x+1) = log base4(20)
------
log4[x(x+1)] = log4(20)
-----
x^2+x-20 = 0
(x+5)(x-4) = 0
Positive solution:
x = 4
===============
Cheers,
Stan H.